let Q( x,y) denote the statement 'x=y+5' , what is the truth value of Q(3,4)


1
Expert's answer
2021-03-05T07:28:00-0500

Q(3,4)



"x=y+5""L.H.S\\x=3\\
R.H.S\\
y+5=4+5\\
=9\\"

since Xis not equal to y+5, the truth value of Q(x,y) is FALSE

Ans: FALSE